OpenStudy (anonymous):

Does the equation represent a direct variation? If so, find the constant of variation. 5x = -3y

OpenStudy (anonymous):

Yes it does but which according to the question is your independent variable. If it is "x", then the variable y varies by \[-\frac{ 5 }{ 3 }\] of x

OpenStudy (anonymous):

x= -3/5*y and y=-5/3*x are the variation constants if you put x= m then y would be -5/3x same for both the variables :)